# Topological characterization of phase transitions and critical edge states in one-dimensional non-Hermitian systems with sublattice symmetry

### v1

@@ III.4 $(\alpha,\alpha^{\prime},\alpha^{\prime\prime})$ chain: critical topological phase transitions
Meanwhile, the $Q(z)$ has for zeros of order one at $z_{\pm}^{{\rm L,R}}$ and a pole of order two at $z=0$. The locations of these zeros and poles versus the GBZ determine the topological invariants of the system following Eqs. (6) and (7).
To proceed, we need to obtain the GBZ from the two middle solutions of the equation $Q(z)=Q(ze^{i\theta})$ for $\theta\in[0,2\pi]$ [26]. As a further simplification, we take $J_{\beta}^{{\rm L}}=J_{\beta}^{{\rm R}}=J_{\beta}\in\mathbb{C}$ with $\beta=\alpha,\alpha+1,\alpha+2$ for the hopping amplitudes and set $J_{\alpha+1}=1$ as the unit of energy. The system Hamiltonian remains non-Hermitian whenever there exists a $\beta\in\{\alpha,\alpha+1,\alpha+2\}$ such that $J_{\beta}\neq J_{\beta}^{*}$. Substituting these conditions into the equation $Q(z)=Q(ze^{i\theta})$, we obtain
$(J_{\alpha+2}+J_{\alpha})ze^{i\theta}+J_{\alpha}J_{\alpha+2}(e^{i\theta}+1)$
$=$ $J_{\alpha}J_{\alpha+2}z^{4}e^{i2\theta}(e^{i\theta}+1)+(J_{\alpha+2}+J_{\alpha})z^{3}e^{i2\theta}.$ (35)
@@
For each given set of parameters $(J_{\alpha},J_{\alpha+2},\theta)$, this equation has four solutions in $z$, with two of them given by $z_{\pm}=\pm e^{-i\theta/2}$. As these solutions have equal magnitudes $|z_{+}|=|z_{-}|=1$, they must be the two middle solutions that determine the GBZ of the system. Therefore, the GBZ radius is given by $r_{0}=|z_{\pm}|=1$, which means that it is identical to the standard BZ. Nevertheless, we will show that there are also topological phases and transitions induced by non-Hermitian effects, which are captured by our theoretical framework.
Figure 9: Topological phase diagram of the NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain. System parameters are $J_{\alpha}^{{\rm L}}=J_{\alpha}^{{\rm R}}=J_{\alpha}=1.5$, $J_{\alpha+1}^{{\rm L}}=J_{\alpha+1}^{{\rm R}}=J_{\alpha+1}=1$ and $J_{\alpha+2}^{{\rm L}}=J_{\alpha+2}^{{\rm R}}=J_{\alpha+2}=J+i\gamma$, where $J,\gamma\in\mathbb{R}$. Each region with a uniform color refers to a gapped phase, with its topological invariant $W$ and zero/pole locations $|z_{\pm}^{{\rm L}}|,|z_{\pm}^{{\rm R}}|$ relative to GBZ shown explicitly therein. The red dashed and solids lines correspond to two sets of topologically distinct phase boundaries with $W=\alpha$ and $W=\alpha+1$, respectively.
@@
In Fig. 9, we show the topological phase diagram of NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain for a representative case. Each continuous region with a uniform color corresponds to a gapped phase, while the red dashed and dotted lines depict topological phase boundaries. In the region surrounded by the red dashed lines, we have $|z_{\pm}^{{\rm L}}|>1$ and $|z_{\pm}^{{\rm R}}|<1$. The winding numbers of $P(z)$ and $Q(z)$ are thus given by $W_{1}=2\alpha$ and $W_{2}=0$, yielding the topological invariant $W=\alpha$ for the corresponding phase. In the region sandwiched between the red dashed and solid lines, we have $|z_{+}^{{\rm L}}|<1$, $|z_{-}^{{\rm L}}|>1$, $|z_{+}^{{\rm R}}|<1$ and $|z_{-}^{{\rm R}}|>1$. The winding numbers of $P(z)$ and $Q(z)$ are thus given by $W_{1}=2\alpha+2$ and $W_{2}=0$, leading to the topological invariant $W=\alpha+1$ for this phase. In the region outside the red solid lines, we have $|z_{\pm}^{{\rm L}}|<1$ and $|z_{\pm}^{{\rm R}}|>1$. The winding numbers of $P(z)$ and $Q(z)$ are then given by $W_{1}=2\alpha+4$ and $W_{2}=0$, generating the topological invariant $W=\alpha+2$ for the associated phase. The red dashed line in Fig. 9 is given by the solution of $|z_{+}^{{\rm L}}|=|z_{-}^{{\rm R}}|=1$. On this dashed line, we have $|z_{-}^{{\rm L}}|>1$ and $|z_{+}^{{\rm R}}|<1$. Therefore, along the phase boundary depicted by the red dashed line, we find the winding numbers of $P(z)$ and $Q(z)$ as $W_{1}=2\alpha+1$ and $W_{2}=-1$, yielding the topological invariant $W=\alpha$. We conclude that the red dashed critical line in Fig. 9 is topologically trivial (nontrivial) if $\alpha=0$ ($\alpha>0$). On the other hand, the red solid line in Fig. 9 is given by the solution of $|z_{-}^{{\rm L}}|=|z_{+}^{{\rm R}}|=1$. On this solid line, we find $|z_{+}^{{\rm L}}|<1$ and $|z_{-}^{{\rm R}}|>1$. Therefore, along the phase boundary depicted by the red solid line, the winding numbers of $P(z)$ and $Q(z)$ are $W_{1}=2\alpha+3$ and $W_{2}=-1$, yielding the topological invariant $W=\alpha+1$. We conclude that the red solid critical line in Fig. 9 is topologically nontrivial for all $\alpha\geq 0$.
@@
We next illustrate our results about the NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain with numerical examples. To be explicit, we focus on the case with $\alpha=0$, set the second neighbor hopping $J_{2}=J+i\gamma$ as the complex parameter, and fix other system parameters as $(J_{0},J_{1},J)=(1.5,1,1)$. The resulting energy spectrum is shown in Fig. 10(a). With the increase of $\gamma$, we observe two phase transitions accompanied by the closing/reopening of bulk spectral gaps (highlighted by the red solid line) at $\gamma\approx 0.65$ and $\gamma\approx 1.75$. After each transition, more eigenmodes appear in the spectrum under the OBC (blue dots) at zero energy. These zero modes are spatially localized, as exemplified by their inverse participation ratios in the spectra reported in Figs. 10(b) and 10(c) for $\gamma=1.2$ and $2.4$, respectively. These zero modes are also found to stay at the two edges of the system, as shown in Figs. 10(d) and 10(f). Further counting reveals that with the raise of $\gamma$, the number of edge zero modes we obtained following the two sequential topological transitions are $N_{0}=2$ and $4$. Meanwhile, according to the phase diagram Fig. 9, the three gapped phases in Fig. 10(a) from left to right have the topological invariants $W=0$, 11 and $2$, successively. The bulk-edge correspondence $N_{0}=2|W|$, as predicted by our theory, is thus verified for all gapped phases of the NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain with $\alpha=0$. We have also checked and confirmed that other choices of $\alpha>0$ generate consistent results.
Figure 11: Spectra and edge states of the NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain along the critical line. We set $\alpha=0,$ $J_{\alpha}^{{\rm L}}=J_{\alpha}^{{\rm R}}=J_{\alpha}=1.5$, $J_{\alpha+1}^{{\rm L}}=J_{\alpha+1}^{{\rm R}}=J_{\alpha+1}=1$, and $J_{\alpha+2}^{{\rm L}}=J_{\alpha+2}^{{\rm R}}=J_{\alpha+2}=J+i\gamma$. In (a) and (b), the values of $J$ and $\gamma$ are taken along the red dashed (solid) line in Fig. 9 with $J\in[0.5,1.5]$ ($\gamma\in(0,1]$) for $\gamma\leq 0$ ($\gamma>0$). (a) and (b) show the real and imaginary parts of the spectrum under the OBC, where the red stars highlight the location of multicritical point at $(J,\gamma)=(1.5,0)$, and the color of each data point gives the inverse participation ratio of the corresponding state. (c) shows the spectrum at $\gamma=0.5$ in (a) and (b) on the complex plane. (a)–(c) share the same color bar. (d) shows the probability distributions of the two critical edge modes at $E=0$ in (c) for a lattice with $100$ unit cells.
Finally, we demonstrate the emergence of critical edge states following phase transitions along non-Hermitian topological phase boundaries in Fig. 11. We focus on a parameter path in Fig. 9 along the dashed critical line with $J\in[0.5,1.5]$ for $\gamma\leq 0$, followed by the solid critical line with $\gamma\in(0,1]$ for $\gamma>0$. These two critical lines, which have been identified as topologically distinct according to our theory, are interconnected at the multicritical point $(J,\gamma)=(1,0)$ in Fig. 9. Interestingly, we observe that by passing through the multicritical point from the dashed to solid phase boundaries, localized eigenmodes are generated at zero energy without undergoing any gap closing and reopening processes in the complex spectrum, as presented in Figs. 11(a) and 11(b). Moreover, the emerging zero modes are indeed coexistent with a gapless bulk and localized spatially around the two ends of the lattice, as shown in Figs. 11(c) and 11(d). These zero modes are thus referred to as non-Hermitian critical edge states, whose topological origin following the “phase transition of phase transition” along the gapless critical line is well described by our theory. Indeed, the phase diagram reported in Fig. 9 predicts $W=0$ and $W=1$ along the dashed and solid phase boundaries. The bulk-edge correspondence $N_{0}=2|W|$ is then verified not only in gapped phases, but also along all gapless phase boundaries of the NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain. Following our definition of topological invariant $W$, we could now arrive at a consistent description of topological phases and transitions in 1D, two-band, sublattice-symmetric non-Hermitian systems. This is regardless of whether the bulk spectrum is gapped or gapless, or whether the underlying Brillouin zone is in Bloch or generalized non-Bloch forms.
@@ IV Summary and Discussion
In this work, we revealed and characterized non-Hermiticity induced critical edge states and topologically nontrivial phase transition points in 1D non-Hermitian systems with sublattice symmetry. By applying the Cauchy’s argument principle to two characteristic functions of a non-Hermitian two-band Hamiltonian, we obtain a pair of winding numbers, whose combination yields an integer-quantized topological invariant $W$. The applicability of this proposed invariant to the overall depiction of gapped topological phases, gapless phase boundaries, topological phase transitions and bulk-edge correspondence is demonstrated by investigating a broad class of non-Hermitian SSH models. Notably, we identify a phase transition induced by non-Hermitian effects along topological phase boundaries, which is accompanied by the quantized jump of our proposed invariant $W$ and the emergence of critical edge zero modes. The origin of these nontrivial gapless topologies could not be consistent described by existing approaches relying on the presence of point or line spectral gaps. One key advantage of our theoretical approach lies in its generality, in the sense that it does not concern whether the bulk bands are separated from edge states of the system by gaps, and whether the underlying band theory is in Bloch or non-Bloch forms. Therefore, we expect our theory to be applicable to the topological characterization of phases and transitions in other more complicated non-Hermitian, two-band models in one dimension with sublattice symmetries.
@@
In future work, it would be interesting to consider the generalization of our theory to systems with more than two bulk bands, in other symmetry classes and beyond one spatial dimension. The extension of our theory to driven systems would also be of great importance for revealing gapless Floquet topology [76, 77, 78] beyond Hermitian limits. Finally, the realization of our NHSSH $(0,1,2)$ chain in quantum simulators like electrical, photonic and acoustic systems may lead to the first experimental observation of critical edge modes and topologically nontrivial critical points in non-Hermitian systems, which deserve more thorough explorations.

### v2

@@ III.3 $(\alpha,\alpha^{\prime},\alpha^{\prime\prime})$ chain: critical topological phase transitions
Meanwhile, the $Q(z)$ has for zeros of order one at $z_{\pm}^{{\rm L,R}}$ and a pole of order two at $z=0$. The locations of these zeros and poles vs the GBZ determine the topological invariants of the system following Eqs. (3) and (4).
@@
To proceed, we need to obtain the GBZ from the two middle solutions of the equation $Q(z)=Q(ze^{i\theta})$ for $\theta\in[0,2\pi]$ [26]. As a further simplification, we take $J_{\beta}^{{\rm L}}=J_{\beta}^{{\rm R}}=J_{\beta}\in\mathbb{C}$ with $\beta=\alpha,\alpha+1,\alpha+2$ for the hopping amplitudes and set $J_{\alpha+1}=1$ as the unit of energy. The system Hamiltonian remains non-Hermitian whenever there exists a $\beta\in\{\alpha,\alpha+1,\alpha+2\}$ such that $J_{\beta}\neq J_{\beta}^{*}$. Substituting these conditions into the equation $Q(z)=Q(ze^{i\theta})$, we obtain $(J_{\alpha+2}+J_{\alpha})ze^{i\theta}+J_{\alpha}J_{\alpha+2}(e^{i\theta}+1)=J_{\alpha}J_{\alpha+2}z^{4}e^{i2\theta}(e^{i\theta}+1)+(J_{\alpha+2}+J_{\alpha})z^{3}e^{i2\theta}$. For each given set of parameters $(J_{\alpha},J_{\alpha+2},\theta)$, this equation has four solutions in $z$, with two of them given by $z_{\pm}=\pm e^{-i\theta/2}$. As these solutions have equal magnitudes $|z_{+}|=|z_{-}|=1$, they must be the two middle solutions that determine the GBZ of the system. Therefore, the GBZ radius is given by $r_{0}=|z_{\pm}|=1$, which means that it is identical to the standard BZ. Nevertheless, we will show that there are still topological phases and transitions induced by non-Hermitian effects, which are captured by our theory.
[…]
@@
In Fig. 7, we show the topological phase diagram of NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain for a representative case. Each region with a uniform color corresponds to a gapped phase, while the red dashed and dotted lines depict topological phase boundaries. In the region surrounded by the red dashed lines, we have $|z_{\pm}^{{\rm L}}|>1$ and $|z_{\pm}^{{\rm R}}|<1$. The winding numbers of $P(z)$ and $Q(z)$ are thus given by $W_{1}=2\alpha$ and $W_{2}=0$, yielding the topological invariant $W=\alpha$ for this phase. In the region sandwiched between the red dashed and solid lines, we have $|z_{+}^{{\rm L}}|<1$, $|z_{-}^{{\rm L}}|>1$, $|z_{+}^{{\rm R}}|<1$ and $|z_{-}^{{\rm R}}|>1$. The winding numbers of $P(z)$ and $Q(z)$ are thus $W_{1}=2\alpha+2$ and $W_{2}=0$, leading to the topological invariant $W=\alpha+1$ for this phase. In the region outside the red solid lines, we have $|z_{\pm}^{{\rm L}}|<1$ and $|z_{\pm}^{{\rm R}}|>1$. The winding numbers of $P(z)$ and $Q(z)$ are then given by $W_{1}=2\alpha+4$ and $W_{2}=0$, generating the topological invariant $W=\alpha+2$ for this phase. The red dashed line in Fig. 7 is given by the solution of $|z_{+}^{{\rm L}}|=|z_{-}^{{\rm R}}|=1$. On this dashed line, we have $|z_{-}^{{\rm L}}|>1$ and $|z_{+}^{{\rm R}}|<1$. Therefore, along the phase boundary depicted by the red dashed line, we find the winding numbers of $P(z)$ and $Q(z)$ as $W_{1}=2\alpha+1$ and $W_{2}=-1$, yielding the topological invariant $W=\alpha$. The red dashed critical line in Fig. 7 is thus topologically trivial (nontrivial) if $\alpha=0$ ($\alpha>0$). On the other hand, the red solid line in Fig. 7 is given by the solution of $|z_{-}^{{\rm L}}|=|z_{+}^{{\rm R}}|=1$. Along this line, we find $|z_{+}^{{\rm L}}|<1$ and $|z_{-}^{{\rm R}}|>1$. Therefore, on the phase boundary depicted by the red solid line, the winding numbers of $P(z)$ and $Q(z)$ are $W_{1}=2\alpha+3$ and $W_{2}=-1$, yielding the topological invariant $W=\alpha+1$. The red solid critical line in Fig. 7 is thus topologically nontrivial for all $\alpha\geq 0$.
@@
We next illustrate our results about the NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain with numerical examples. To be explicit, we focus on the case with $\alpha=0$, set the second neighbor hopping $J_{2}=J+i\gamma$ as the complex parameter, and fix other system parameters as $(J_{0},J_{1},J)=(1.5,1,1)$. The resulting energy spectrum is shown in Fig. 8(a). With the increase of $\gamma$, we observe two phase transitions accompanied by the closing/reopening of bulk spectral gaps (highlighted by the red solid line) at $\gamma\approx 0.65$ and $\gamma\approx 1.75$. After each transition, more eigenmodes appear in the spectrum under the OBC (blue dots) at zero energy. These zero modes are spatially localized, as exemplified by their inverse participation ratios in the spectra reported in Figs. 8(b) and 8(c) for $\gamma=1.2$ and $2.4$, respectively. These zero modes are also found to stay at the two edges of the system, as shown in Figs. 8(d) and 8(e). Further counting reveals that with the raise of $\gamma$, the number of edge zero modes we obtained following the two sequential topological transitions are $N_{0}=2$ and $4$. Meanwhile, according to the phase diagram Fig. 7, the three gapped phases in Fig. 8(a) from left to right have the topological invariants $W=0$, 11 and $2$, successively. The bulk-edge correspondence $N_{0}=2|W|$, as predicted by our theory, is thus verified for all gapped phases of the NHSSH $(\alpha,\alpha+1,\alpha+2)$ chain with $\alpha=0$. We have also checked and confirmed that other choices of $\alpha>0$ generate consistent results.
[…]
@@ IV Summary and Discussion
In this work, we revealed and characterized non-Hermiticity induced critical edge states and topologically nontrivial phase transition points in 1D non-Hermitian systems with sublattice symmetry. By applying the Cauchy’s argument principle to two characteristic functions of a non-Hermitian two-band Hamiltonian, we obtain a pair of winding numbers, whose combination yields an integer-quantized topological invariant $W$. The applicability of this proposed invariant to the overall depiction of gapped topological phases, gapless phase boundaries, topological phase transitions and bulk-edge correspondence is demonstrated by investigating a broad class of non-Hermitian SSH models. Notably, we identify a phase transition induced by non-Hermitian effects along topological phase boundaries, which is accompanied by the quantized jump of our proposed invariant $W$ and the emergence of critical edge zero modes. The origin of these nontrivial gapless topology could not be consistently described by existing approaches relying on the presence of point or line spectral gaps. One key advantage of our theoretical approach lies in its generality, in the sense that it does not concern whether the bulk bands are separated from edge states of the system by gaps, and whether the underlying band theory is in Bloch or non-Bloch forms. Therefore, we expect our theory to be applicable to the topological characterization of phases and transitions in other more complicated non-Hermitian, two-band models in one dimension with sublattice symmetry.
One possible way of interpreting the underling physical mechanism of our found gSPT phases may rest on the concept of symmetry-enriched quantum criticality [57]. We illustrate this point here with an explicit example. The global symmetry that protects the topologically nontrivial critical points of our system is the sublattice symmetry, which reads ${\hat{\Gamma}}=\sum_{n}({\hat{a}}_{n}^{\dagger}{\hat{a}}_{n}-{\hat{b}}_{n}^{\dagger}{\hat{b}}_{n})$ for NHSSH $\alpha$ chains in the lattice representation. This symmetry enforces an edge zero mode to occupy only one type of sublattice A or B. Therefore, under the OBC, we may identify two sublattice-resolved fermionic parity symmetries at the edges, which are given by ${\hat{P}}_{1,{\rm A}}=1-2{\hat{a}}_{1}^{\dagger}{\hat{a}}_{1}$ and ${\hat{P}}_{N,{\rm B}}=1-2{\hat{b}}_{N}^{\dagger}{\hat{b}}_{N}$, where $N$ is the total number of unit cells. For the NHSSH $(0,1)$ chain, the symmetrized Hamiltonian under the OBC reads ${\hat{H}}^{\prime}_{01}={\hat{H}}^{\prime}_{01,{\rm Ledge}}+{\hat{H}}^{\prime}_{01,{\rm bulk}}+{\hat{H}}^{\prime}_{01,{\rm Redge}}$, where ${\hat{H}}^{\prime}_{01,{\rm Ledge}}=\varrho^{-1}(J+\gamma){\hat{a}}_{1}^{\dagger}{\hat{b}}_{1}+\varrho(J-\gamma){\hat{b}}_{1}^{\dagger}{\hat{a}}_{1}$, ${\hat{H}}^{\prime}_{01,{\rm Redge}}=\varrho^{-1}(J+\gamma){\hat{a}}_{N}^{\dagger}{\hat{b}}_{N}+\varrho(J-\gamma){\hat{b}}_{N}^{\dagger}{\hat{a}}_{N}$, and the ${\hat{H}}^{\prime}_{01,{\rm bulk}}$ is given by Eq. (36) with the summation taken over n=1,,N1n=1,...,N-1. At the critical point of the NHSSH $(0,1)$ chain, the ${\hat{H}}^{\prime}_{01}$ transforms under the ${\hat{P}}_{1,{\rm A}}$ and ${\hat{P}}_{N,{\rm B}}$ as ${\hat{P}}_{1,{\rm A}}{\hat{H}}^{\prime}_{01}{\hat{P}}_{1,{\rm A}}=-{\hat{H}}^{\prime}_{01,{\rm Ledge}}+{\hat{H}}^{\prime}_{01,{\rm bulk}}+{\hat{H}}^{\prime}_{01,{\rm Redge}}$ and ${\hat{P}}_{N,{\rm B}}{\hat{H}}^{\prime}_{01}{\hat{P}}_{N,{\rm B}}={\hat{H}}^{\prime}_{01,{\rm Ledge}}+{\hat{H}}^{\prime}_{01,{\rm bulk}}-{\hat{H}}^{\prime}_{01,{\rm Redge}}$. Therefore, the parity symmetries ${\hat{P}}_{1,{\rm A}}$ and ${\hat{P}}_{N,{\rm B}}$ are both broken at the edges. Their only common eigenstate with the system Hamiltonian is the trivial vacuum state $|\emptyset\rangle$, with ${\hat{P}}_{1,{\rm A}}|\emptyset\rangle=|\emptyset\rangle$, ${\hat{P}}_{N,{\rm B}}|\emptyset\rangle=|\emptyset\rangle$, and ${\hat{H}}^{\prime}_{01}|\emptyset\rangle=0$. On the other hand, the symmetrized Hamiltonian of NHSSH $(1,2)$ chain under the OBC is given by Eq. (43), with the first (second) summation taken over n=1,,N1n=1,...,N-1 (n=1,,N2n=1,...,N-2). At its critical point, we have ${\hat{P}}_{1,{\rm A}}{\hat{H}}^{\prime}_{12}{\hat{P}}_{1,{\rm A}}={\hat{P}}_{N,{\rm B}}{\hat{H}}^{\prime}_{12}{\hat{P}}_{N,{\rm B}}={\hat{H}}^{\prime}_{12}$. Therefore, both the parity symmetries ${\hat{P}}_{1,{\rm A}}$ and ${\hat{P}}_{N,{\rm B}}$ are preserved at the edges of NHSSH $(1,2)$ chain. Each of these symmetries share a nontrivial eigenstate with ${\hat{H}}^{\prime}_{12}$ at one edge, i.e., ${\hat{P}}_{1,{\rm A}}({\hat{a}}_{1}^{\dagger}|\emptyset\rangle)=-({\hat{a}}_{1}^{\dagger}|\emptyset\rangle)$, ${\hat{P}}_{N,{\rm B}}({\hat{b}}_{N}^{\dagger}|\emptyset\rangle)=-({\hat{b}}_{N}^{\dagger}|\emptyset\rangle)$, ${\hat{H}}^{\prime}_{12}({\hat{a}}_{1}^{\dagger}|\emptyset\rangle)=0({\hat{a}}_{1}^{\dagger}|\emptyset\rangle)$ and ${\hat{H}}^{\prime}_{12}({\hat{b}}_{N}^{\dagger}|\emptyset\rangle)=0({\hat{b}}_{N}^{\dagger}|\emptyset\rangle)$. The symmetries ${\hat{P}}_{1,{\rm A}}$ and ${\hat{P}}_{N,{\rm B}}$ thus allow two eigenmodes to appear at the two edges of the NHSSH $(1,2)$ chain even when the system is critical. The topological degeneracy of these edge modes at zero energy is further protected by the sublattice symmetry. In comparison to the trivial critical point of the NHSSH $(0,1)$ chain, one may regard the ${\hat{P}}_{1,{\rm A}}$ and ${\hat{P}}_{N,{\rm B}}$ as extra symmetries that enable the presence of critical edge modes at a topologically nontrivial gapless point. While the explicit form of edge fermionic parity operators could be model-dependent, their existence may serve as a general mechanism to understand the appearance of gSPT phases in other 1D non-Hermitian systems with sublattice symmetry.
As a second remark, we emphasize that it is the sublattice symmetry $^{\hat{\Gamma}}$ that protects the edge modes in the gapped phases and along the gapless phase boundaries in our non-Hermitian system. The $^{\hat{\Gamma}}$ is anti-commute with the system Hamiltonian $^{\hat{H}}$, which means that it enforces energy eigenstates to appear in pairs with opposite signs. To see this, we note that if $|\psi\rangle$ is an eigenstate of $^{\hat{H}}$ with energy $E$, then ${\hat{\Gamma}}{\hat{H}}|\psi\rangle={\hat{\Gamma}}{\hat{H}}{\hat{\Gamma}}{\hat{\Gamma}}|\psi\rangle=-{\hat{H}}{\hat{\Gamma}}|\psi\rangle={\hat{\Gamma}}E|\psi\rangle=E{\hat{\Gamma}}|\psi\rangle$, so that $H({\hat{\Gamma}}|\psi\rangle)=-E({\hat{\Gamma}}|\psi\rangle)$. Therefore, ${\hat{\Gamma}}|\psi\rangle$ must be an eigenstate of $^{\hat{H}}$ with energy E-E. If the eigenenergy $E=0$, the states $|\psi\rangle$ and ${\hat{\Gamma}}|\psi\rangle$ would become degenerate. Therefore, degenerate zero modes of a sublattice-symmetric Hamiltonian must come in pairs. On the other hand, it was known that the topological phase of a 1D, free-fermion lattice model with sublattice symmetry is characterized by an integer-valued topological invariant, which is the $W$ in our theory. In a phase with $W\neq 0$ (either gapped or gapless), we will find $N_{0}=2|W|$ zero-energy eigenmodes at the two edges of the 1D chain in the thermodynamic limit. Each of these edge zero modes has a degenerate partner at zero energy, with their topological degeneracy being protected by the sublattice symmetry. If this phase is gapless, there could also be bulk band touching point at zero energy, whose degeneracy is again enforced by the sublattice symmetry.
Finally, since our theory is applicable to both Hermitian and non-Hermitian cases, the topology it describes does not directly concern whether the critical point is related to a Hermitian-type normal degeneracy or a non-Hermitian exceptional point (EP). In our case studies of NHSSH $(0,1)$ and $(1,2)$ chains, there could be an infinite-order bulk EP within bulk bands at $J=\gamma=\sqrt{2}$ in Figs. 3(a) and 5(a), where the bulk bands become flat at $|E|=1$. However, in parameter regions at and close to these EPs, the two bulk bands are well gapped at $E=0$, as can be seen from Figs. 3(a) and 5(a). Therefore, the topology at system edges remains insensitive to these bulk EPs. Overall, even though the EP forms a unique type of non-Hermitian degeneracy, its impact does not show up directly in the topological states we characterized.
@@
In future work, it would be interesting to consider the generalization of our theory to systems with more than two bulk bands [98, 99, 100, 101], in other symmetry classes, beyond one spatial dimension, or with impurities [102, 103, 104]. Topologically nontrivial criticality and critical edge states that are originated solely from an exceptional-point degeneracy would be an intriguing direction to explore. The extension of our theory to driven systems would also be of great importance for revealing gapless Floquet topology [80, 81, 82] beyond Hermitian limits. On application side, the topologically protected non-Hermitian criticality may offer alternative routes to the quantum-enhanced sensing in lattice systems [105]. Finally, the realization of our NHSSH $(0,1,2)$ chain in quantum simulators like electrical, photonic and acoustic systems [106, 107, 108, 109, 110] may lead to the first experimental observation of critical edge modes and topologically nontrivial critical points in non-Hermitian systems, which deserve more thorough explorations.